OpenStudy (anonymous):

Can someone check my answer for: Simplify 11√8 +√12 - 5√2 I checked my answer on Wolfram and it says the answer is 17√2 + 12√3 But this is how I got my answer: 11√8 = √2 * 4 = √2 * √4 = 2√4 = 2√2 6√12 = √3 * 4 = √3 * √4 = 3√2 - 5√2 cannot be simplified anymore 2 + 3 - 5 = 0 √2 + √2 + √2 = √2 Combine everything: √2 Have I done something wrong, or do I have the right answer?

OpenStudy (anonymous):



OpenStudy (anonymous):

hope the helped

OpenStudy (ghazi):

\[11 \sqrt 8 + \sqrt 12-5 \sqrt2= 11*2 \sqrt2+ 2 \sqrt3-5 \sqrt2= 22 \sqrt2-5 \sqrt2+2 \sqrt3\]

OpenStudy (ghazi):

\[17 \sqrt2 +2 \sqrt3\] should be the answer

OpenStudy (anonymous):

how come the 6 was taken out of 6√12?

OpenStudy (ghazi):

\[\sqrt{12}= \sqrt {2*2*3}= 2 \sqrt 3\]

OpenStudy (anonymous):

Oh, I just relooked my question! Sorry. I forgot the 6.

OpenStudy (ghazi):

its okay :)